Rearrange V(t)/Vo = e^-t/RC to solve for C
Answer #1

Here it can be written as

V (t)/ Vo = e -t/RC
Since we need to solve for C so first we have to get rid of e

To remove e we take log both sides as log(V(t)/Vo)=log(e-t/RC)
So, logV(t) - logVo = -t/RC

[When you divide two numbers, you subtract their logs]

Now C=-t/[(logv(t)-logvo)(R)]